# Confidential Databases Without Cryptographic Mappings

### 19 Mar 2026 v1

@@ 7.2 Sensitivity Analysis (RQ2)
The performance gain of Fedb depends on both workload characteristics and system configuration. To characterize the impact of these factors, we use Sysbench oracle2021sysbench to issue transactional microbenchmarks with configurable read/write ratios and access patterns. The benchmark runs on a synthetic dataset consisting of 32 tables with 1M rows each, occupying approximately 7 GB, 10 GB, and 12 GB in plaintext PostgreSQL, Fedb, and HEDB, respectively. To control for other variables, we set the total available memory equal to the machine’s physical memory, except in the memory-sensitivity experiments. When experimenting with memory sensitivity, we use cgroups to adjust the total available memory (the sum of the integrity and privacy zones) to match different sizes of the dataset. Each experiment runs a 120-second warm-up followed by a 600-second measurement, using 16 client connections by default.
@@ 7.3 Miscellaneous (RQ3)
Performance on Intel x86. Our evaluation is based on ARM commodity hardware, as it relies on ARM TEE features Li2021TwinVisor. We also measure operation costs on an Intel Xeon x86 platform that supports AES-NI hardware acceleration. The results on x86 show that en/decryption remains an order of magnitude slower than put/get operations, as detailed in Table 4.
@@ 8 Security Analysis
Consistent with prior modern CDBs Vinayagamurthy2019StealthDB; Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; Li2023HEDB, we consider two adversary classes: active adversaries can arbitrarily tamper with storage and modify untrusted memory, and passive adversaries (honest-but-curious) do not modify data but attempt to infer sensitive information through observation.
Defense against active attacks. Fedb ensures that sensitive data remains confidential and fresh. First, TEE hardware provides memory encryption and isolation, preventing adversaries from accessing plaintext data during computation. Second, when data is evicted to untrusted storage, Fedb encrypts it using authenticated encryption at the block level with a counter-based freshness mechanism arm-rollback-protection, which can detect storage corruption and rollback attacks Angel2023Nimble; chu2026rollbaccine.
Fedb ensures faithful query execution through two mechanisms. First, remote attestation allows users to verify legitimate Fedb code is running in the TEEs before uploading data. Second, following HEDB Li2023HEDB, Fedb places both the DBMS (in the integrity zone) and expression operators (in the privacy zone) within trusted domains, with secure channels between TEEs that prevent message injection, replay, man-in-the-middle and interface attacks.
Defense against passive attacks. The primary question is whether an adversary observing FIDs can learn anything about the corresponding plaintext secrets. We prove that Fedb provides indistinguishability. The observation is that FID allocation is stateful but plaintext-independent. The allocator maintains an internal counter (tracking the next available FID) and assigns FIDs deterministically from the counter, independent of the plaintext . After prior Put() calls, any two plaintexts and submitted as the -th item receive the exact same FID. Thus, the adversary’s view of the FID sequence is identical for and . Fedb’s crypto-free mappings provide confidentiality comparable to semantic security goldwasser1982probabilistic: no information about plaintexts is revealed (beyond leakage from other channels, discussed later).
@@
Leakage analysis. While Fedb protects plaintext contents, practical query functionality requires revealing certain metadata to the untrusted infrastructure. We characterize Fedb’s leakage profile using the definition of -security Poddar2016Arx; Vinayagamurthy2019StealthDB, which models security as a leakage function that precisely specifies what information the adversary learns. We define Fedb’s as follows:
@@
Comparison results: the boolean outcomes (e.g., “”) necessary for indexing and filtering, a common tradeoff to many enterprise CDBs Antonopoulos2020AEv2; Guo2021GaussDB; Wang2022Operon; Li2023HEDB.
@@
Side-channel considerations. Fedb’s stateful privacy zone retains plaintext data in TEE memory for extended periods (until eviction to storage). This residency could increase the attack window for digital side-channels (e.g., cache timing or memory access patterns). Modern CDBs already face similar risks, as encryption keys reside permanently in TEE memory—key compromise would breach all sensitive data. Existing mitigation for key protection (e.g., strict CPU resource isolation zhou2023coreslicing; castes2024coregapping) can equally safeguard plaintext data in TEEs. To protect against controlled-channel attacks xu2015controlledchannel, Fedb utilizes ARM Secure EL2 armv8-manual that fully isolates TEEs and their page tables from untrusted privileged software. Importantly, Fedb’s design is generalizable to other TEEs offering equivalent guarantees, including Intel TDX intel-tdx, AMD SEV-SNP amd-sev, IBM PEF Hunt2021PEF, etc. Physical attacks that infer DRAM access patterns remain prohibitively expensive Lee2020OffChip.
@@ 9 Related Work
Confidential databases (CDBs). To address rising privacy concerns, CDBs on untrusted clouds have long been a goal Maheshwari2000TDB. Systems like TrustedDB Bajaj2011TrustedDB, EnclaveDB Priebe2018EnclaveDB, DBStore Ribeiro2018DBStore, and SecuDB yang2024secudb shield an entire DBMS in TEEs for strong security with little maintainability. Modern CDBs, like AlwaysEncrypted Antonopoulos2020AEv2, GaussDB Guo2021GaussDB, Operon Wang2022Operon, and HEDB Li2023HEDB prioritize maintainability, yet often incur severe performance overhead. Motivated by the growing industrial adoption of this latter class of systems, Fedb aims to improve their efficiency while preserving their advantages.
Other CDB systems take different approaches: fully cryptographic databases Popa2011CryptDB; Poddar2016Arx; Tu2013Monomi; bian2023he3db; ren2022heda; bian2025engorgio; zhang2024arcedb; zhao2025hermes require no trusted hardware, and oblivious databases mishra2018oblix; crooks2018obladi; zhu2025compass hide access patterns. These lines of work are orthogonal to ours.
Indirection-protection separation. Separating indirection from protection is a classic systems principle lampson2020hints. Capability systems dennis1980capability; levy2014capability pioneered the use of unforgeable tokens (capabilities) as references to resources protected by OS isolation. Similarly, OSes have long used opaque handles ritchie1974unix; accetta1986mach: file descriptors in Unix are integers mapped by the kernel to internal objects. Fedb shares this insight: data-independent FIDs serve as capabilities protected by TEE isolation, removing cryptographic overhead on every reference.
Applicability beyond CDBs. Distributed confidential analytics systems, like Spark-based Opaque zheng2017opaque and Laputa Kim2025Laputa, Hadoop-based Civet Tsai2020Civet and VC3 schuster2015vc3, encrypt data during cross-node shuffles, incurring costly cryptographic operations at TEE boundaries. These systems may benefit from the abstraction of crypto-free mappings for better efficiency.
TEE Optimization. Several techniques have been proposed to optimize TEE-based applications, such as context switchless Antonopoulos2020AEv2; Arnautov2016SCONE; Arasu2013Cipherbase; weisse2017hotcalls, userspace I/O bailleu2019speicher; Thalheim2021rktio, and others kim2019shieldstore; Li2023HEDB.
@@ Acknowledgments
This work was sponsored by CCF-Huawei Populus Grove Fund. Mingyu Li and Haibo Chen are the corresponding authors.

### 23 Mar 2026 v2

@@ 7.2 Sensitivity Analysis (RQ2)
The performance gain of Fedb depends on both workload characteristics and system configuration. To characterize the impact of these factors, we use Sysbench [48] to issue transactional microbenchmarks with configurable read/write ratios and access patterns. The benchmark runs on a synthetic dataset consisting of 32 tables with 1M rows each, occupying approximately 7 GB, 10 GB, and 12 GB in plaintext PostgreSQL, Fedb, and HEDB, respectively. To control for other variables, we set the total available memory equal to the machine’s physical memory, except in the memory-sensitivity experiments. When experimenting with memory sensitivity, we use cgroups to adjust the total available memory (the sum of the integrity and privacy zones) to match different sizes of the dataset. Each experiment runs a 120-second warm-up followed by a 600-second measurement, using 16 client connections by default.
@@ 7.3 Miscellaneous (RQ3)
Performance on Intel x86. Our evaluation is based on ARM commodity hardware, as it relies on ARM TEE features [24]. We also measure operation costs on an Intel Xeon x86 platform that supports AES-NI hardware acceleration. The results on x86 show that en/decryption remains an order of magnitude slower than put/get operations, as detailed in Table 4.
@@ 8 Security Analysis
Consistent with prior modern CDBs [10, 12, 14, 13, 15], we consider two adversary classes: active adversaries can arbitrarily tamper with storage and modify untrusted memory, and passive adversaries (honest-but-curious) do not modify data but attempt to infer sensitive information through observation.
Defense against active attacks. Fedb ensures that sensitive data remains confidential and fresh. First, TEE hardware provides memory encryption and isolation, preventing adversaries from accessing plaintext data during computation. Second, when data is evicted to untrusted storage, Fedb encrypts it using authenticated encryption at the block level with a counter-based freshness mechanism [49], which can detect storage corruption and rollback attacks [50, 51].
Fedb ensures faithful query execution through two mechanisms. First, remote attestation allows users to verify legitimate Fedb code is running in the TEEs before uploading data. Second, following HEDB [15], Fedb places both the DBMS (in the integrity zone) and expression operators (in the privacy zone) within trusted domains, with secure channels between TEEs that prevent message injection, replay, man-in-the-middle and interface attacks.
Defense against passive attacks. The primary question is whether an adversary observing FIDs can learn anything about the corresponding plaintext secrets. We prove that Fedb provides indistinguishability. The observation is that FID allocation is stateful but plaintext-independent. The allocator maintains an internal counter (tracking the next available FID) and assigns FIDs deterministically from the counter, independent of the plaintext . After prior Put() calls, any two plaintexts and submitted as the -th item receive the exact same FID. Thus, the adversary’s view of the FID sequence is identical for and . Fedb’s crypto-free mappings provide confidentiality comparable to semantic security [52]: no information about plaintexts is revealed (beyond leakage from other channels, discussed later).
@@
Leakage analysis. While Fedb protects plaintext contents, practical query functionality requires revealing certain metadata to the untrusted infrastructure. We characterize Fedb’s leakage profile using the definition of -security [53, 10], which models security as a leakage function that precisely specifies what information the adversary learns. We define Fedb’s as follows:
@@
Side-channel considerations. Fedb’s stateful privacy zone retains plaintext data in TEE memory for extended periods (until eviction to storage). This residency could increase the attack window for digital side-channels (e.g., cache timing or memory access patterns). Modern CDBs already face similar risks, as encryption keys reside permanently in TEE memory—key compromise would breach all sensitive data. Existing mitigation for key protection (e.g., strict CPU resource isolation [54, 55]) can equally safeguard plaintext data in TEEs. To protect against controlled-channel attacks [56], Fedb utilizes ARM Secure EL2 [45] that fully isolates TEEs and their page tables from untrusted privileged software. Importantly, Fedb’s design is generalizable to other TEEs offering equivalent guarantees, including Intel TDX [19], AMD SEV-SNP [17], IBM PEF [25], etc. Physical attacks that infer DRAM access patterns remain prohibitively expensive [57].
@@ 9 Related Work
Confidential databases (CDBs). To address rising privacy concerns, CDBs on untrusted clouds have long been a goal [58]. Systems like TrustedDB [7], EnclaveDB [9], DBStore [32], and SecuDB [16] shield an entire DBMS in TEEs for strong security with little maintainability. Modern CDBs, like AlwaysEncrypted [12], GaussDB [14], Operon [13], and HEDB [15] prioritize maintainability, yet often incur severe performance overhead. Motivated by the growing industrial adoption of this latter class of systems, Fedb aims to improve their efficiency while preserving their advantages.
Other CDB systems take different approaches: fully cryptographic databases [59, 53, 60, 61, 62, 63, 64, 65] require no trusted hardware, and oblivious databases [66, 67, 68] hide access patterns. These lines of work are orthogonal to ours.
Indirection-protection separation. Separating indirection from protection is a classic systems principle [35]. Capability systems [36, 37] pioneered the use of unforgeable tokens (capabilities) as references to resources protected by OS isolation. Similarly, OSes have long used opaque handles [38, 39]: file descriptors in Unix are integers mapped by the kernel to internal objects. Fedb shares this insight: data-independent FIDs serve as capabilities protected by TEE isolation, removing cryptographic overhead on every reference.
Applicability beyond CDBs. Distributed confidential analytics systems, like Spark-based Opaque [69] and Laputa [70], Hadoop-based Civet [71] and VC3 [72], encrypt data during cross-node shuffles, incurring costly cryptographic operations at TEE boundaries. These systems may benefit from the abstraction of crypto-free mappings for better efficiency.
@@ Acknowledgments
This work was sponsored by CCF-Huawei Populus Grove Fund. Mingyu Li is the corresponding author.